solve the system
4x+9y=5
2x-6y=-1

Answer by checkley77(12844)

4x+9y=5
2x-6y=-1 multiply this equation by -2 & add.
-4x+12y=2
4x+9y=5
------------------
21y=7
y=7/21
y=1/3 answer.
4x+9*1/3=5
4x+9/3=5
4x+3=5
4x=5-3
4x=2
x=2/4
x=1/2 answer.
Proof:
2*1/2-6*1/3=-1
1-6/3=-1
1-2=-1
-1=-1